INVARIANTS OF ORTHOGONAL G-MODULES FROM THE CHARACTER TABLE

GABRIELE NEBE

Abstract. Character-theoretic methods using Clifford algebras are developed to describe the quadratic forms on a vector space V that are invariant under a finite subgroup G of GL(V) such that the order of the commutator factor group of G is odd.